OpenStudy (anonymous):

Find the value of x and y in the following equation. 7+4y=-3x 7x+108=9y

OpenStudy (anonymous):

do you know how to start a problem like this? You first just want to get one of the variable alone so you can plug what it equals into the other equation.

OpenStudy (anonymous):

Yeah I know that process

OpenStudy (anonymous):

ok so what did you get stuck on?

OpenStudy (anonymous):

how would you change the problems so that they match each other?

OpenStudy (anonymous):

what u want to do is get ur equation to look like this: x =..... then u just plug what x equals into the other equation.

OpenStudy (anonymous):

ok so far I have: -3x+4y=7 and 7x+9y=108. What number would I multiply one of the equations so they cancel?

OpenStudy (anonymous):

ok so take 7x + 9y = 108 then subtract 7x 9y = 108 -7x then divide by 9 y = 7/9x + 12 now you can plug this for y in the other equation.